OpenStudy (anonymous):

Solve by using the substitution method. 8x+4y=-44 -2x+y=21 What is the solution of the system? HELP PLEASE

OpenStudy (anonymous):

does anyone know the answer? I need it ASAP

OpenStudy (anonymous):

You want to cancel one of the terms out to solve for one of the terms: 8x+4y=-44 4(-2x+y)=21(4) so... 8x+4y=-44 -8x+4y=84 the 8x's cancel out, leaving... 4y=-44 (+)4y=84 equals 8y=40 y=5

OpenStudy (anonymous):

Now substitute y=5 into either of the equations and solve for x

OpenStudy (anonymous):

8x+4(5)=-44; 8x+20=-44; 8x=-64; x=-8 so, x=-8 and y=5

OpenStudy (anonymous):

so if written as an order pair the solution would be (-8,5)?

OpenStudy (anonymous):

That's the substitution method - you want to cancel out a term to solve for the other term, then substitute back in to solve for the other term.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Got it , thanks!

OpenStudy (anonymous):

No prob.

OpenStudy (anonymous):

I have some other question i need help with . I'll post them in a new question